OpenStudy (lawlesscrob):

The distance you travel while hiking is a function of how fast you hike and how long you hike at this rate. You usually maintain a speed of three miles per hour while hiking. Write an equation that represents this function, letting t represent the independent variable, let h represent the function, and h(t) the dependent variable. Then, determine h(4). a. h(t)(3t/4); 48 miles b. t(h)=3h; 12 miles c. h(t)=3t; 12 miles d. h(t)=(3t/4); 3 miles

OpenStudy (lapinblanc):

C. A way to remember this is to think about how you cancel out to find your distance. Your speed is in miles per hour (represented as miles/hour). You are trying to find miles and they give your speed: 3miles/hour So in order to end up with just miles, you multiply by time or (t). h(t)=h*t=ht so h(t)=3*t, h(4)=3t input 4 into t h(4)=3(4)=12miles Slightly more detailed explanation: You know time is the independent variable because no matter how fast you go, time will not change no matter how fast you hike. So in h(t) t is time, and its unit is hr. speed is miles/hr, you multiply to cancel out the hr and just get minutes.

OpenStudy (lawlesscrob):

Thank you @lapinblanc

OpenStudy (lapinblanc):

You are very welcome.